OpenStudy (tasia):

Evaluate: |2.8 - k| + 3.4 if k = 8.1

OpenStudy (anonymous):

|2.8 - 8.1| + 3.4 = 8.7

OpenStudy (anonymous):

|2.8 - 8.1| + 3.4 = | -5.3 | + 3.4 = 5.3 + 3.4 Add this to get the answer...